Find the wavelength for a typical am radio wave with frequency 1000KHz


1
Expert's answer
2021-07-02T09:39:51-0400

The wavelength of an electromagnetic wave is related with frequency by equation

"\lambda=\frac{c}{f}=\rm\frac{3*10^8\: m/s}{1000*10^3\: Hz}=300\: m"